A perticle moving in semicircle of radius 6 m with constant speed of 12 m/s .calculate average velocities
1
Expert's answer
2019-08-21T08:55:37-0400

Displacement of the particle = diameter of the semicircle "d = 2*6 = 12 m."

Particle motion time "t=L\/v," where "L" is length of the semicircle, "v" is speed of the particle along the semicircle.

Average velocity "v_{av}=d/t=d/L\cdot v=2\cdot 6 /(\pi\cdot 6)\cdot 12 = 24/\pi" m/s

Answer: "24/\pi" m/s.